OpenStudy (anonymous):

Expand ln 4x^2y Please show work if needed.:)

OpenStudy (lalaly):

\[\ln (a \times b)= \ln (a) +\ln(b)\]so \[\ln(4x^2 \times y) = \ln (4x^2)+\ln(y)\]and \[\ln (a^b)=bln(a)\]so\[\ln(4x^2 \times y) = 2 \ln (4x)+\ln(y)\]

OpenStudy (anonymous):

Thank you can you also tell me if the final answer for \[\ln 12/5\] is \[\ln 12-\ln 5 ?\]

OpenStudy (lalaly):

if its \[\ln(\frac{12}{5})\] then yeah you got it right:D

OpenStudy (anonymous):

Thank you! I like the roses (in your picture) :D

OpenStudy (lalaly):

hehe Thanks your sweet:D